# Equivariant Ehrhart theory, commutative algebra and invariant triangulations of polytopes

### v2

@@ 2.4. Equivariant Ehrhart theory
As in Example 2.3, assume that $d$ is an even positive integer, and let $r=\frac{d}{2}+1$. Let $a_{1},\ldots,a_{r}$ be positive integers with $\gcd(a_{1},\ldots,a_{r})=1$. Let $e_{1},\ldots,e_{r},f_{1},\ldots,f_{r}$ be a basis of $\mathbb{Z}^{d+2}$. Let $\bar{e}_{i}$ and $\bar{f_{i}}$ denote the images of $e_{i}$ and $f_{i}$ respectively in the lattice $L:=\mathbb{Z}^{d+2}/\mathbb{Z}(\sum_{i}a_{i}(e_{i}-f_{i}))$. Consider the group homomorphism $\psi:L\to\mathbb{Z}$ satisfying $\psi(\bar{e}_{i})=\psi(\bar{f}_{i})=1$ for all $1\leq i\leq r$, and the corresponding linear map $\psi_{\mathbb{R}}:\mathbb{R}^{d+1}\to\mathbb{R}$. As in Example 2.4, $M^{\prime}=\psi^{-1}(1)$ is an affine lattice in the real affine space $W=\psi_{\mathbb{R}}^{-1}(1)$. Let $P=\operatorname{Conv}\left\{{\bar{e}_{1},\ldots,\bar{e}_{r},\bar{f}_{1},\ldots,\bar{f}_{r}}\right\}\subset W$. Then $P$ is a lattice polytope with respect to $M^{\prime}$ with $22r=d+$ 2 vertices and normalized volume $\operatorname{Vol}(P)=\sum_{i}a_{i}$. Consider an affine representation $\rho:G\to\operatorname{Aff}(M^{\prime})$ such that $P$ is $G$-invariant. Suppose there exists an element $g$ in $G$ such that $g\cdot\{\bar{e}_{1},\ldots,\bar{e}_{r}\}=\{\bar{f}_{1},\ldots,\bar{f}_{r}\}$. After possibly relabelling coordinates, we may assume that there exists a partition $I_{1},\ldots,I_{s}$ of $\{1,\ldots,r\}$ such that the $\langle g\rangle$-orbits of the vertices of $P$ are precisely $\{\bar{e_{i}},\bar{f_{i}}:i\in I_{k}\}$ for $1\leq k\leq s$. By [Sta11a, Lemma 5.4], the fixed locus $P^{g}$ is the rational simplex with vertices $\{\frac{1}{2|I_{k}|}\sum_{i\in I_{k}}(\bar{e}_{i}+\bar{f}_{i}):1\leq k\leq s\}$. By Example 2.3,
@@
We deduce that $h^{*}(P,\rho;t)(g)$ is a polynomial if and only if |$|\{a_{i}:a_{i}\textrm{ odd }\}|$ is odd. The latter condition holds if and only if $\operatorname{Vol}(P)=\sum_{i}a_{i}$ is odd.
@@
[Sta11a, Corollary 6.6] Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensonal lattice polytope. For any $g$ in $G$, we have the following equality of rational functions:
@@
This follows from Remark 2.7. Explicitly, if $g_{i}$ in $G_{i}$ acts on $(\widetilde{M}_{i})_{\mathbb{C}}$ with eigenvalues $\{\lambda_{i,1},\ldots,\lambda_{i,d_{i}}\}$ for $i\in\{1,2\}$, then $(g_{1},g_{2})$ acts on $(\widetilde{M}_{1})_{\mathbb{C}}\oplus(\widetilde{M}_{2})_{\mathbb{C}}$ with eigenvalues $\{\lambda_{1,1},\ldots,\lambda_{1,d_{1}},\lambda_{2,1},\ldots,\lambda_{2,d_{2}}\}$. ∎
@@
With the notation above, let $G_{1}=G$, $\rho_{1}=\rho$ and $P_{1}=P$. Let $G_{2}=\{1\}$ act on $P_{2}=M_{2}=\{0\}$. Then the free join $P_{1}*P_{2}$ is the pyramid $\operatorname{Pyr}(P)$ of $P$. The pyramid $\operatorname{Pyr}(P)$ may be identified with the the convex hull of $P\times\{1\}$ and the origin in $\widetilde{M}_{\mathbb{R}}$ with $G$ acting via $\rho:G\to\operatorname{GL}(\widetilde{M})$. In this case, Lemma 2.22 implies that $(1-t)\operatorname{Ehr}(\operatorname{Pyr}(P),\rho;t)=\operatorname{Ehr}(P,\rho;t)$ and $h^{*}(\operatorname{Pyr}(P),\rho;t)=h^{*}(P,\rho;t)$ (see [Sta11a, Remark 6.2]).
@@ 2.5. The equivariant $h^{*}_{N}$-series
We consider the setup of Remark 2.12. As in Example 2.4, consider the map $\psi:\mathbb{Z}^{d+1}\to\mathbb{Z}$, $\psi(u_{1},\ldots,u_{d+1})=\sum_{i=1}^{d+1}u_{i}$ and corresponding map $\psi_{\mathbb{R}}:\mathbb{R}^{d+1}\to\mathbb{R}$. Then $M^{\prime}=\psi^{-1}(\binom{d+2}{2})$ is an affine lattice in the real affine space $W=\psi_{\mathbb{R}}^{-1}(\binom{d+2}{2})$. The symmetric group $G=\operatorname{Sym}_{d+1}$ acts on $M^{\prime}$ by permuting coordinates. Let $\rho:G\to\operatorname{Aff}(M^{\prime})$ denote the corresponding affine representation of $G$. Let $P$ be the convex hull of the orbit $G\cdot(1,2,\ldots,d+1)\subset W$. Then $P$ is the $(d+1)$-permutahedron. The equivariant Ehrhart theory of $P$ was studied in [ASVM20] and [ASVM21]. In particular, explicit combinatorial formulas for $L(P,\rho;m)$ and $h^{*}(P,\rho;t)(g)$ were given in [ASVM20, Theorem 1.1] and [ASVM20, Proposition 5.1] respectively. In particular, they showed that if $g$ has cycle type $(\ell_{1},\ldots,\ell_{r})$, then $h^{*}(P,\rho;t)(g)$ has a pole at $t=-1$ if and only if $h^{*}(P,\rho;t)(g)$ is not a polynomial if and only if the number $e$ of even parts of the cycle type of $g$ satisfies $10<e<r-$ 1 [ASVM20, Lemma 5.2]. They deduced that $h^{*}(P,\rho;t)$ is a polynomial if and only if $h^{*}(P,\rho;t)$ is effective if and only if $d\leq 2$ [ASVM20, Proposition 5.3-5.4]. Since $\det(I-\widetilde{M}_{\mathbb{C}}^{\prime}t)(g)=\prod_{i}(1-t^{\lambda_{i}})$, we compute:
@@
It follows that if $h^{*}(P,\rho;t)(g)$ is not a polynomial and $N$ is odd, then $h^{*}_{N}(P,\rho;t)(g)$ is not a polynomial. When $N$ is even, we will see in Example 3.7 that $P$ admits a $G$-invariant triangulation with vertices in $\frac{1}{N}M^{\prime}$ and hence Corollary 1.6 implies that $h^{*}_{N}(P,\rho;t)$ is a polynomial. We conclude that $h^{*}_{N}(P,\rho;t)(g)$ is not a polynomial if and only if $N$ is odd and the number $e$ of even parts of the cycle type of $g$ satisfies $10<e<r-$ 1.
@@
[Sta11a, Corollary 6.6] Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensonal lattice polytope. Fix a positive integer $N$. For any $g$ in $G$, we have the following equality of rational functions:
@@
Recall that $\rho_{N}:G\to\operatorname{Aff}(\frac{1}{N}M)\subset\operatorname{GL}(\frac{1}{N}\widetilde{M})$ denotes the corresponding affine representation of $G$ acting on $\frac{1}{N}M$. Recall from Example 2.23 that we may consider the induced action of $G$ on the pyramid $\operatorname{Pyr}(P)$ of $P$.
@@ 3.1. Polyhedral subdivisions of polytopes
We say that $\mathcal{S}$ is $G$-invariant if $g\cdot F\in\mathcal{S}$ for all $g$ in $G$ and $F\in\mathcal{S}$. Elements of $\mathcal{S}$ are called faces of $\mathcal{S}$, and the 00-dimensional and maximal faces of $\mathcal{S}$ are called vertices and facets respectively. We say that $\mathcal{S}$ is a lattice or rational polytopal complex if the vertices of $\mathcal{S}$ lie in $M$ or $M_{\mathbb{Q}}$ respectively. The support $\operatorname{Supp}(\mathcal{S})$ of $\mathcal{S}$ is the union of the faces $\{F_{i}:i\in I\}$ in $M_{\mathbb{R}}$. We say that $\mathcal{S}$ is a polyhedral subdivision of its support. In particular, a polyhedral subdivision of $P$ is a polytopal complex in $M_{\mathbb{R}}$ with support $P$. A triangulation of $P$ is a polyhedral subdivision of $P$ such that each facet is a simplex i.e. has precisely $d+1$ vertices.
@@ 3.2. Constructing invariant triangulations
Our main tool is the following lemma. Below we consider the empty set to be a simplex of dimension $-1$.
@@
Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M_{\mathbb{R}})$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensonal polytope. Let $u_{1},\ldots,u_{r}$ be a collection of points in $P$ and let $\mathcal{S}$ be a $G$-invariant regular polyhedral subdivision of $P$. Consider the subcomplex $\mathcal{T}(i):=\{F\in\mathcal{S}:F\cap(\cup_{1\leq j\leq i}G\cdot u_{j})=\emptyset\}$ of $\mathcal{S}$ for all $0\leq i\leq r$. Assume that the following properties are satisfied:
@@
Let $\mathcal{S}(0)=\mathcal{S}$ and let $\mathcal{S}(i)=\mathcal{S}(i-1)_{G\cdot u_{i}}$ for $1\leq i\leq r$. By Remark 3.4, $\mathcal{S}(r)$ is a $G$-invariant regular polyhedral subdivision of $P$. It remains to show that $\mathcal{S}(r)$ is a triangulation. Recall that given polytopes $Q,Q^{\prime},Q^{\prime\prime}$ in $M_{\mathbb{R}}$ with $Q=\operatorname{Conv}\left\{{Q^{\prime},Q^{\prime\prime}}\right\}$ we say that $Q$ is the join of $Q^{\prime}$ and $Q^{\prime\prime}$ and write $Q=Q^{\prime}*Q^{\prime\prime}$ if $\dim Q=\dim Q^{\prime}+\dim Q^{\prime\prime}+1$. We claim that every facet $F$ of $\mathcal{S}(i)$ has the form $F=F^{\prime}*F^{\prime\prime}$ where $F^{\prime}\in\mathcal{T}(i)$ and $F^{\prime\prime}$ is a simplex with vertices contained in $\cup_{1\leq j\leq i}G\cdot u_{j}$. Since all nonempty faces of $\mathcal{T}(r)$ are simplices by assumption, it follows that $\mathcal{S}(r)$ is a triangulation. It remains to prove the claim. The claim holds when $i=0$ since $\mathcal{S}(0)=\mathcal{T}(0)=\mathcal{S}$ and we may write $F=F^{\prime}*F^{\prime\prime}$ with $F^{\prime}=F$ and $F^{\prime\prime}=\emptyset$ for any facet $F$ of $\mathcal{S}(0)$. Assume that $1\leq i\leq r$. By assumption, $u_{i}\in\operatorname{Supp}(\mathcal{T}(i-1))$ and since $\mathcal{T}(i-1)$ is $G$-invariant, the orbit $G\cdot u_{i}$ lies in $\operatorname{Supp}(\mathcal{T}(i-1))$. Let $F$ be a facet of $\mathcal{S}(i-1)$. It is enough to establish the claim for all facets of $\mathcal{S}(i)|_{F}=\mathcal{S}(i-1)_{G\cdot u_{i}}|_{F}$. By induction, we can write $F=F^{\prime}*F^{\prime\prime}$ where $F^{\prime}\in\mathcal{T}(i-1)$ and $F^{\prime\prime}$ is a simplex with vertices contained in $\cup_{1\leq j\leq i-1}G\cdot u_{j}$. Then $F^{\prime}$ is the intersection of $F$ with $\operatorname{Supp}(\mathcal{T}(i-1))$. From our assumptions, $|F\cap(G\cdot u_{i})|=|F^{\prime}\cap(G\cdot u_{i})|\leq 1$. If $F\cap(G\cdot u_{i})=\emptyset$, then $F^{\prime}\in\mathcal{T}(i)$ and $F\in\mathcal{S}(i)$ by Lemma 3.5, and the claim follows. It remains to consider the case when $F\cap(G\cdot u_{i})=\{w\}$ is a point. Let $H$ be a facet of $F$ that doesn’t contain $w$. Then $H\cap(G\cdot u_{i})=\emptyset$. Since $F=F^{\prime}*F^{\prime\prime}$ and $w\in F^{\prime}$, we have $H=H^{\prime}*F^{\prime\prime}$ for some facet $H^{\prime}$ of $F^{\prime}$. Since $H^{\prime}\subset F^{\prime}\in\mathcal{T}(i-1)$, we have $H^{\prime}\in\mathcal{T}(i)$. By Lemma 3.5, the polytopes of the form $\operatorname{Conv}\left\{{H,w}\right\}$ are precisely the facets of $\mathcal{S}(i)|_{F}$. Then $\operatorname{Conv}\left\{{H,w}\right\}=H^{\prime}*\operatorname{Conv}\left\{{F^{\prime\prime},w}\right\}$, where $H^{\prime}\in\mathcal{T}(i)$ and $\operatorname{Conv}\left\{{F^{\prime\prime},w}\right\}$ is a simplex with vertices contained in $\cup_{1\leq j\leq i}G\cdot u_{j}$. This completes the proof of the claim. ∎
@@
We continue with the setup of Example 2.26. That is, $\psi:\mathbb{Z}^{d+1}\to\mathbb{Z}$, $\psi(u_{1},\ldots,u_{d+1})=\sum_{i=1}^{d+1}u_{i}$, and $M^{\prime}=\psi^{-1}(\binom{d+2}{2})$ is an affine lattice in the real affine space $W=\psi_{\mathbb{R}}^{-1}(\binom{d+2}{2})$. The symmetric group $G=\operatorname{Sym}_{d+1}$ acts on the $(d+1)$-permutahedron $P\subset W$. The nonempty faces of $P$ are in order-preserving bijection with the set of ordered partitions of $\{1,\ldots,d+1\}$ (see, for example, [Zie95, Example 0.10]). Explicitly, consider an ordered partition $I=(I_{1},\ldots,I_{s})$ of $\{1,\ldots,d+1\}$ and the corresponding face $F_{I}$ of $P$. Then $\dim F_{I}=d+1-s$ and the vertices of $F_{I}$ are precisely all $v=(v_{1},\ldots,v_{d+1})$ such that we have an equality of unordered sets $\{v_{i}:i\in I_{1}\cup\cdots\cup I_{j}\}=\{1,2,\ldots,\sum_{k\leq j}|I_{k}|\}$ for all $1\leq j\leq s$. Then the average of the vertices of $F_{I}$ is the element $u_{I}=(u_{I,1},\ldots,u_{I,d+1})$ defined by
@@
Let $\mathcal{S}$ be the trivial subdivision of $P$. Our goal is to inductively construct a sequence $u_{1},\ldots,u_{r}$ in $P\cap\frac{1}{N}M$ for some positive integer $r$ and apply Lemma 3.6. As in Lemma 3.6, we will consider the subcomplex $\mathcal{T}(i):=\{F\subset P:F\cap(\cup_{1\leq j\leq i}G\cdot u_{j})=\emptyset\}$ of $P$ for all $0\leq i\leq r$. Let $u_{1}$ to be a $G$-fixed point in $P\cap\frac{1}{N}M$. This exists since for any $u\in P\cap M$, the average $\frac{1}{|G|}\sum_{g\in G}g\cdot u\in P\cap\frac{1}{N}M$ is $G$-fixed. Suppose that we have constructed $u_{1},\ldots,u_{i-1}\in P\cap\frac{1}{N}M$ for some positive integer $i>1$. If $\mathcal{T}(i-1)$ is empty, then set $r=i-1$. Otherwise, let $u$ be a lattice point in $\operatorname{Supp}(\mathcal{T}(i-1))$. Let $m=\max_{Q\in\mathcal{T}(i-1)}|Q\cap(G\cdot u)|\in\mathbb{Z}_{>0}$. Fix $Q$ in $\mathcal{T}(i-1)$ such that $|Q\cap(G\cdot u)|=m$. After possibly replacing $Q$ by a face of $Q$ we may also assume that $Q$ is the smallest face of $P$ containing $Q\cap(G\cdot u)$. Let $G_{Q}$ denote the stabilizer of $Q$ and let $u_{i}$ be a $G_{Q}$-fixed point in the relative interior $Q^{\circ}$ of $Q$.

### v3

@@ 2.4. Equivariant Ehrhart theory
As in Example 2.3, assume that $d$ is an even positive integer, and let $r=\frac{d}{2}+1$. Let $a_{1},\ldots,a_{r}$ be positive integers with $\gcd(a_{1},\ldots,a_{r})=1$. Let $e_{1},\ldots,e_{r},f_{1},\ldots,f_{r}$ be a basis of $\mathbb{Z}^{d+2}$. Let $\bar{e}_{i}$ and $\bar{f_{i}}$ denote the images of $e_{i}$ and $f_{i}$ respectively in the lattice $L:=\mathbb{Z}^{d+2}/\mathbb{Z}(\sum_{i}a_{i}(e_{i}-f_{i}))$. Consider the group homomorphism $\psi:L\to\mathbb{Z}$ satisfying $\psi(\bar{e}_{i})=\psi(\bar{f}_{i})=1$ for all $1\leq i\leq r$, and the corresponding linear map $\psi_{\mathbb{R}}:\mathbb{R}^{d+1}\to\mathbb{R}$. As in Example 2.4, $M^{\prime}=\psi^{-1}(1)$ is an affine lattice in the real affine space $W=\psi_{\mathbb{R}}^{-1}(1)$. Let $P=\operatorname{Conv}\left\{{\bar{e}_{1},\ldots,\bar{e}_{r},\bar{f}_{1},\ldots,\bar{f}_{r}}\right\}\subset W$. Then $P$ is a lattice polytope with respect to $M^{\prime}$ with $2r=d+2$ vertices and normalized volume $\operatorname{Vol}(P)=\sum_{i}a_{i}$. Consider an affine representation $\rho:G\to\operatorname{Aff}(M^{\prime})$ such that $P$ is $G$-invariant. Suppose there exists an element $g$ in $G$ such that $g\cdot\{\bar{e}_{1},\ldots,\bar{e}_{r}\}=\{\bar{f}_{1},\ldots,\bar{f}_{r}\}$. After possibly relabelling coordinates, we may assume that there exists a partition $I_{1},\ldots,I_{s}$ of $\{1,\ldots,r\}$ such that the $\langle g\rangle$-orbits of the vertices of $P$ are precisely $\{\bar{e_{i}},\bar{f_{i}}:i\in I_{k}\}$ for $1\leq k\leq s$. By [Sta11a, Lemma 5.4], the fixed locus $P^{g}$ is the rational simplex with vertices $\{\frac{1}{2|I_{k}|}\sum_{i\in I_{k}}(\bar{e}_{i}+\bar{f}_{i}):1\leq k\leq s\}$. By Example 2.3,
@@
We deduce that $h^{*}(P,\rho;t)(g)$ is a polynomial if and only if $|\{a_{i}:a_{i}\textrm{ odd }\}|$ is odd. The latter condition holds if and only if $\operatorname{Vol}(P)=\sum_{i}a_{i}$ is odd.
@@
[Sta11a, Corollary 6.6] Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensional lattice polytope. For any $g$ in $G$, we have the following equality of rational functions:
@@
This follows from Remark 2.7. Explicitly, if $g_{i}$ in $G_{i}$ acts on $(\widetilde{M}_{i})_{\mathbb{C}}$ with eigenvalues $\{\lambda_{i,1},\ldots,\lambda_{i,d_{i}}\}$ for $i\in\{1,2\}$, then $(g_{1},g_{2})$ acts on $(\widetilde{M}_{1})_{\mathbb{C}}\oplus(\widetilde{M}_{2})_{\mathbb{C}}$ with eigenvalues $\{\lambda_{1,1},\ldots,\lambda_{1,d_{1}},\lambda_{2,1},\ldots,\lambda_{2,d_{2}}\}$.
@@
[…]
@@ 2.5. The equivariant $h^{*}_{N}$-series
We consider the setup of Remark 2.12. As in Example 2.4, consider the map $\psi:\mathbb{Z}^{d+1}\to\mathbb{Z}$, $\psi(u_{1},\ldots,u_{d+1})=\sum_{i=1}^{d+1}u_{i}$ and corresponding map $\psi_{\mathbb{R}}:\mathbb{R}^{d+1}\to\mathbb{R}$. Then $M^{\prime}=\psi^{-1}(\binom{d+2}{2})$ is an affine lattice in the real affine space $W=\psi_{\mathbb{R}}^{-1}(\binom{d+2}{2})$. The symmetric group $G=\operatorname{Sym}_{d+1}$ acts on $M^{\prime}$ by permuting coordinates. Let $\rho:G\to\operatorname{Aff}(M^{\prime})$ denote the corresponding affine representation of $G$. Let $P$ be the convex hull of the orbit $G\cdot(1,2,\ldots,d+1)\subset W$. Then $P$ is the $(d+1)$-permutahedron. The equivariant Ehrhart theory of $P$ was studied in [ASVM20] and [ASVM21]. In particular, explicit combinatorial formulas for $L(P,\rho;m)$ and $h^{*}(P,\rho;t)(g)$ were given in [ASVM20, Theorem 1.1] and [ASVM20, Proposition 5.1] respectively. In particular, they showed that if $g$ has cycle type $(\ell_{1},\ldots,\ell_{r})$, then $h^{*}(P,\rho;t)(g)$ has a pole at t=1t=-1 if and only if $h^{*}(P,\rho;t)(g)$ is not a polynomial if and only if the number $e$ of even parts of the cycle type of $g$ satisfies 0<e<r10<e<r-1 [ASVM20, Lemma 5.2]. They deduced that $h^{*}(P,\rho;t)$ is a polynomial if and only if $h^{*}(P,\rho;t)$ is effective if and only if $d\leq 2$ [ASVM20, Proposition 5.3-5.4]. Since $\det(I-\widetilde{M}_{\mathbb{C}}^{\prime}t)(g)=\prod_{i}(1-t^{\lambda_{i}})$, we compute:
@@
It follows that if $h^{*}(P,\rho;t)(g)$ is not a polynomial and $N$ is odd, then $h^{*}_{N}(P,\rho;t)(g)$ is not a polynomial. When $N$ is even, we will see in Example 3.7 that $P$ admits a $G$-invariant triangulation with vertices in $\frac{1}{N}M^{\prime}$ and hence Corollary 1.6 implies that $h^{*}_{N}(P,\rho;t)$ is a polynomial. We conclude that $h^{*}_{N}(P,\rho;t)(g)$ is not a polynomial if and only if $N$ is odd and the number $e$ of even parts of the cycle type of $g$ satisfies 0<e<r10<e<r-1.
@@
[Sta11a, Corollary 6.6] Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensional lattice polytope. Fix a positive integer $N$. For any $g$ in $G$, we have the following equality of rational functions:
@@
[…]
@@ 3.1. Polyhedral subdivisions of polytopes
We say that $\mathcal{S}$ is $G$-invariant if $g\cdot F\in\mathcal{S}$ for all $g$ in $G$ and $F\in\mathcal{S}$. Elements of $\mathcal{S}$ are called faces of $\mathcal{S}$, and the 0-dimensional and maximal faces of $\mathcal{S}$ are called vertices and facets respectively. We say that $\mathcal{S}$ is a lattice or rational polytopal complex if the vertices of $\mathcal{S}$ lie in $M$ or $M_{\mathbb{Q}}$ respectively. The support $\operatorname{Supp}(\mathcal{S})$ of $\mathcal{S}$ is the union of the faces $\{F_{i}:i\in I\}$ in $M_{\mathbb{R}}$. We say that $\mathcal{S}$ is a polyhedral subdivision of its support. In particular, a polyhedral subdivision of $P$ is a polytopal complex in $M_{\mathbb{R}}$ with support $P$. A triangulation of $P$ is a polyhedral subdivision of $P$ such that each facet is a simplex i.e. has precisely $d+1$ vertices.
@@ 3.2. Constructing invariant triangulations
Our main tool is the following lemma. Below we consider the empty set to be a simplex of dimension 1-1.
@@
Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M_{\mathbb{R}})$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensional polytope. Let $u_{1},\ldots,u_{r}$ be a collection of points in $P$ and let $\mathcal{S}$ be a $G$-invariant regular polyhedral subdivision of $P$. Consider the subcomplex $\mathcal{T}(i):=\{F\in\mathcal{S}:F\cap(\cup_{1\leq j\leq i}G\cdot u_{j})=\emptyset\}$ of $\mathcal{S}$ for all $0\leq i\leq r$. Assume that the following properties are satisfied:
@@
Let $\mathcal{S}(0)=\mathcal{S}$ and let $\mathcal{S}(i)=\mathcal{S}(i-1)_{G\cdot u_{i}}$ for $1\leq i\leq r$. By Remark 3.4, $\mathcal{S}(r)$ is a $G$-invariant regular polyhedral subdivision of $P$. It remains to show that $\mathcal{S}(r)$ is a triangulation.
Recall that given polytopes $Q,Q^{\prime},Q^{\prime\prime}$ in $M_{\mathbb{R}}$ with $Q=\operatorname{Conv}\left\{{Q^{\prime},Q^{\prime\prime}}\right\}$ we say that $Q$ is the join of $Q^{\prime}$ and $Q^{\prime\prime}$ and write $Q=Q^{\prime}*Q^{\prime\prime}$ if $\dim Q=\dim Q^{\prime}+\dim Q^{\prime\prime}+1$. We claim that every facet $F$ of $\mathcal{S}(i)$ has the form $F=F^{\prime}*F^{\prime\prime}$ where $F^{\prime}\in\mathcal{T}(i)$ and $F^{\prime\prime}$ is a simplex with vertices contained in $\cup_{1\leq j\leq i}G\cdot u_{j}$. Since all nonempty faces of $\mathcal{T}(r)$ are simplices by assumption, it follows that $\mathcal{S}(r)$ is a triangulation.
It remains to prove the claim. The claim holds when $i=0$ since $\mathcal{S}(0)=\mathcal{T}(0)=\mathcal{S}$ and we may write $F=F^{\prime}*F^{\prime\prime}$ with $F^{\prime}=F$ and $F^{\prime\prime}=\emptyset$ for any facet $F$ of $\mathcal{S}(0)$. Assume that $1\leq i\leq r$. By assumption, $u_{i}\in\operatorname{Supp}(\mathcal{T}(i-1))$ and since $\mathcal{T}(i-1)$ is $G$-invariant, the orbit $G\cdot u_{i}$ lies in $\operatorname{Supp}(\mathcal{T}(i-1))$. Let $F$ be a facet of $\mathcal{S}(i-1)$. It is enough to establish the claim for all facets of $\mathcal{S}(i)|_{F}=\mathcal{S}(i-1)_{G\cdot u_{i}}|_{F}$. By induction, we can write $F=F^{\prime}*F^{\prime\prime}$ where $F^{\prime}\in\mathcal{T}(i-1)$ and $F^{\prime\prime}$ is a simplex with vertices contained in $\cup_{1\leq j\leq i-1}G\cdot u_{j}$. Then $F^{\prime}$ is the intersection of $F$ with $\operatorname{Supp}(\mathcal{T}(i-1))$. From our assumptions, $|F\cap(G\cdot u_{i})|=|F^{\prime}\cap(G\cdot u_{i})|\leq 1$. If $F\cap(G\cdot u_{i})=\emptyset$, then $F^{\prime}\in\mathcal{T}(i)$ and $F\in\mathcal{S}(i)$ by Lemma 3.5, and the claim follows.
It remains to consider the case when $F\cap(G\cdot u_{i})=\{w\}$ is a point. Let $H$ be a facet of $F$ that doesn’t contain $w$. Then $H\cap(G\cdot u_{i})=\emptyset$. Since $F=F^{\prime}*F^{\prime\prime}$ and $w\in F^{\prime}$, we have $H=H^{\prime}*F^{\prime\prime}$ for some facet $H^{\prime}$ of $F^{\prime}$. Since $H^{\prime}\subset F^{\prime}\in\mathcal{T}(i-1)$, we have $H^{\prime}\in\mathcal{T}(i)$. By Lemma 3.5, the polytopes of the form $\operatorname{Conv}\left\{{H,w}\right\}$ are precisely the facets of $\mathcal{S}(i)|_{F}$. Then $\operatorname{Conv}\left\{{H,w}\right\}=H^{\prime}*\operatorname{Conv}\left\{{F^{\prime\prime},w}\right\}$, where $H^{\prime}\in\mathcal{T}(i)$ and $\operatorname{Conv}\left\{{F^{\prime\prime},w}\right\}$ is a simplex with vertices contained in $\cup_{1\leq j\leq i}G\cdot u_{j}$. This completes the proof of the claim. ∎
@@
We continue with the setup of Example 2.26. That is, $\psi:\mathbb{Z}^{d+1}\to\mathbb{Z}$, $\psi(u_{1},\ldots,u_{d+1})=\sum_{i=1}^{d+1}u_{i}$, and $M^{\prime}=\psi^{-1}(\binom{d+2}{2})$ is an affine lattice in the real affine space $W=\psi_{\mathbb{R}}^{-1}(\binom{d+2}{2})$. The symmetric group $G=\operatorname{Sym}_{d+1}$ acts on the $(d+1)$-permutahedron $P\subset W$.
The nonempty faces of $P$ are in order-preserving bijection with the set of ordered partitions of $\{1,\ldots,d+1\}$ (see, for example, [Zie95, Example 0.10]). Explicitly, consider an ordered partition $I=(I_{1},\ldots,I_{s})$ of $\{1,\ldots,d+1\}$ and the corresponding face $F_{I}$ of $P$. Then $\dim F_{I}=d+1-s$ and the vertices of $F_{I}$ are precisely all $v=(v_{1},\ldots,v_{d+1})$ such that we have an equality of unordered sets $\{v_{i}:i\in I_{1}\cup\cdots\cup I_{j}\}=\{1,2,\ldots,\sum_{k\leq j}|I_{k}|\}$ for all $1\leq j\leq s$. Then the average of the vertices of $F_{I}$ is the element $u_{I}=(u_{I,1},\ldots,u_{I,d+1})$ defined by
@@
Let $\mathcal{S}$ be the trivial subdivision of $P$. Our goal is to inductively construct a sequence $u_{1},\ldots,u_{r}$ in $P\cap\frac{1}{N}M$ for some positive integer $r$ and apply Lemma 3.6. As in Lemma 3.6, we will consider the subcomplex $\mathcal{T}(i):=\{F\subset P:F\cap(\cup_{1\leq j\leq i}G\cdot u_{j})=\emptyset\}$ of $P$ for all $0\leq i\leq r$.
Let $u_{1}$ to be a $G$-fixed point in $P\cap\frac{1}{N}M$. This exists since for any $u\in P\cap M$, the average $\frac{1}{|G|}\sum_{g\in G}g\cdot u\in P\cap\frac{1}{N}M$ is $G$-fixed. Suppose that we have constructed $u_{1},\ldots,u_{i-1}\in P\cap\frac{1}{N}M$ for some positive integer $i>1$. If $\mathcal{T}(i-1)$ is empty, then set r=i1r=i-1. Otherwise, let $u$ be a lattice point in $\operatorname{Supp}(\mathcal{T}(i-1))$. Let $m=\max_{Q\in\mathcal{T}(i-1)}|Q\cap(G\cdot u)|\in\mathbb{Z}_{>0}$. Fix $Q$ in $\mathcal{T}(i-1)$ such that $|Q\cap(G\cdot u)|=m$. After possibly replacing $Q$ by a face of $Q$ we may also assume that $Q$ is the smallest face of $P$ containing $Q\cap(G\cdot u)$. Let $G_{Q}$ denote the stabilizer of $Q$ and let $u_{i}$ be a $G_{Q}$-fixed point in the relative interior $Q^{\circ}$ of $Q$.